OpenStudy (anonymous):

The cost of a telephone call is 75 cents plus 25 cents times the number of minutes. Write an algebraic expression that models the cost of a telephone call that lasts t minutes.

OpenStudy (camerondoherty):

c = cost c = 75 + 25t

OpenStudy (anonymous):

lol when was your book written, 1976?

OpenStudy (camerondoherty):

It is c = 75 + 25t because c = cost and the initial price is 75 plus 25 per minute So if you talked for 10 minutes it would be c = 75 + 25(10) which would equal $3.25

OpenStudy (anonymous):

Thanks